# Second Language (Arabic) Acquisition of LLMs via Progressive Vocabulary Expansion

### v1

@@ 4.2 Training details
In refining our methodology for the LLaMA2 model’s vocabulary expansion to enhance its handling of Arabic, we not only identified and integrated 12,800 new Arabic subwords using the I-BPE method but also adjusted the language content ratio at each of the 16 training stages  [^3] In principle, a stageless solution could be employed, allowing the addition of one token after another without the need to define the boundaries between stages. However, for the sake of simplifying the implementation, particularly in terms of data preparation, we have opted for a staged approach where we make the number of stages $N=16$.. Each stage involved training with 30B tokens, totaling 480B tokens across all stages. Both English and Arabic data were used at each stage, with the proportion of these languages determined using a cosine annealing schedule. To ensure robust inference capabilities, we included code and mathematical data, maintaining a consistent 5% at every stage, see details in Appendix D. Following the expansion of the vocabulary through the aforementioned stages, to further enhance the model’s performance, we continued training on an additional 20B data based on the expanded vocabulary.
In this paper, we continue pre-training on LLaMA2 models, which have 7 billion (7B) and 13 billion (13B) parameters, using a computational framework composed of 2,368 GPUs. We employ a model parallelism of 2 and a pipeline parallelism of 4. Optimization was carried out using the AdamW optimizer, with a context length of 4,096 tokens for each model. At the start of every training stage, we reintroduced a cosine learning rate scheduler with an initial rate of 1e-5 and decreased to 2e-6, ensuring a gradual adaptation through a 15% warm-up period at the beginning of each stage. Gradient accumulation was set at 8, achieving a total batch size of 4,736 and enabling the processing of approximately 0.019 billion tokens per batch.
@@ 5.1 Experimental settings
Aspect Benchmark Language (+ translation) Size Evaluation Types Metrics
Knowledge Ability RACE Lai et al. (2017) EN 4.9K Multiple-choice Questions Accuracy
MMLU Hendrycks et al. (2021a) EN (+AR) 14K Multiple-choice Questions Accuracy
Arabic Cultural and Value Alignment ACVA-all Huang et al. (2023b) AR 9K Yes/No binary Questions F1-score
ACVA-clean AR 2.48K Yes/No binary Questions F1-score
Commonsense Reasoning BoolQ Clark et al. (2019) EN (+AR) 3.27K Yes/No binary Questions Accuracy
Table 1: Overview of Evaluation benchmarks
@@
As shown in Table 1, we employ four popular benchmarks aimed at assessing world knowledge: (1) MMLU (Measuring Massive Multitask Language Understanding) - This dataset is designed to measure the knowledge acquired during pretraining. For this benchmark, we employ both the original English version from Hendrycks et al. (2021b) and the Arabic version proposed by Huang et al. (2023b), ensuring comprehensive coverage. (2) RACE (Reading Comprehension from Examinations) - A large-scale reading comprehension dataset designed to evaluate the educational knowledge of the models. (3) EXAMS (Multi-subject High School Examinations Dataset for Cross-lingual and Multilingual Question Answering) - Different from the previous benchmarks, EXAMS provides a diverse range of subjects for evaluation. (4) ArabicMMLU - Similar to the global MMLU, this dataset is specifically tailored for original Arabic LLMs, encompassing various countries and subjects. Additionally, evaluating Arabic cultural and value alignment is crucial. To assess this, we utilize ACVA-all and ACVA-clean for localization testing. To comprehensively evaluate model performance on inference and reasoning ability, we translate two commonsense reasoning benchmarks of varying difficulty: BoolQ and ARC-Challenge (ARC-C).
To ensure a fair comparison of candidate models, we adhere to the settings established for each benchmark separately. Furthermore, for translated benchmarks, we utilize the generation approach evaluation method as outlined in Huang et al. (2023b). Specifically, we employed GPT-3.5-Turbo-1106 to translate datasets from English to Arabic for benchmarks that were not originally in Arabic.
@@
To compare LLMs trained or available in Arabic, we have selected several prominent Arabic LLMs or multilingual LLMs as baselines for comparison: (1) AceGPT-[7B,13B] Huang et al. (2023b): This set includes fully fine-tuned generative text models based on LLaMA2, specifically customized for the Arabic language domain. (2) Mistral-7B-Instruct-v0.2 Jiang et al. (2023): The fine-tuned model achieves a balance between performance and efficiency. (3) Jais-[13B,30B] Sengupta et al. (2023): A pre-trained bilingual large language model designed for both Arabic and English. (4) Bloom-[7B]: A multilingual language model extensively trained on diverse textual data, allowing it to produce fluent text in 46 languages and 13 programming languages. (5) LLaMA2-[7B,13B]: A popular and competitive baseline model in the general domain. (6) OpenAI GPT: This includes GPT4 and ChatGPT, closed-source LLMs also strong at multilingual tasks.
@@ 5.2 Evaluation Results
In our study, the performance of base models was assessed on two Arabic-specific MMLU datasets: Arabic MMLU translate Huang et al. (2023b) and ArabicMMLU Koto et al. (2024). The left side of Table 10 details the models’ accuracies on the Arabic MMLU translate dataset within a few-shot setting. It is evident from the data that the AraLLaMA-7B-base and AraLLaMA-13B-base models exhibit superior accuracy rates compared to models of similar scale. Notably, the AraLLaMA-13B-base model outperforms the Jais-30B model, which has a significantly larger parameter count.
@@
Models Arabic-trans MMLU Huang et al. (2023b) ArabicMMLU Koto et al. (2024) Total
STEM Human- ities Social Sciences Others Avg. STEM Social Sciences Human- ities Arabic Language Other Avg. Avg.
@@
ChatGPT 3.5 Turbo 43.38 44.12 55.57 53.21 49.07 53.8 57.0 57.5 57.6 63.8 57.7 53.39
Table 2: Evaluation of base models. We adopt a few-shot setting on Arabic-translated MMLU Huang et al. (2023b) and a zero-shot setting with option logit probability in ArabicMMLU Koto et al. (2024). Numbers with the best performance are in bold in 7B and 13B groups.
@@
MMLU (trans) MMLU Koto et al. (2024) EXAMS ACVA clean ACVA all BoolQ (trans) ARC-C (trans) Avg. BoolQ RACE Avg. Avg.
@@
Table 3 presents the comprehensive evaluation results across various benchmarks for the candidate models, spanning from Arabic to English. Overall, AraLLaMA outperforms all baseline models in the Arabic language tasks. Particularly noteworthy is its proficiency in knowledge-related evaluations such as Arabic-translated MMLU and EXAMS, surpassing other models by at least 1.3%. This highlights the model’s expertise in addressing Arabic knowledge-related questions. Additionally, AraLLaMA demonstrates strong performance in tasks related to Arabic culture and value alignment. In terms of commonsense reasoning, AraLLaMA exhibits notable skills in tasks such as the translated versions of BoolQ and ARC-Challenge, showcasing its reasoning capabilities in Arabic. Beyond Arabic benchmarks, we also investigated the English proficiency of the models to determine whether specialization in one language affects performance in the other. The results indicate that the model maintains its English proficiency and displays robustness in multilingual assessments. It is noteworthy that the lower accuracy of the Jais is attributed to its refusal to answer for unknown reasons.
In a comprehensive evaluation of the ACVA dataset aimed at gauging the understanding of Arabic cultural nuances under a zero-shot setting, our AraLLaMA models showcased unparalleled performance. The AraLLaMA-13B-chat, in particular, stood out with exceptional Average F1 scores of 76.37% and 76.90% in “all set" and "clean Set" categories, respectively, even outperforming the renowned ChatGPT 3.5 Turbo in the "All set" category. This performance not only highlights the AraLLaMA models’ superior grasp of Arabic culture but also establishes them as leading figures among open-source models in this nuanced domain. Compared to other top-tier open-source contenders, including the Jais-30B-chat variants, the AraLLaMA-13B-chat model’s superior results. The instruction-following tests can be found in Appendix H.
@@ 6 More Analysis
Model STEM Social Sciences Humanities Arabic Language Other Avg.
@@ 6.1 Ablation Study on Progressive Vocabulary Expansion
To further demonstrate the effectiveness of progressive vocabulary expansion in downstream task adaptation, we conduct continuous pre-training on a 1B-parameter TinyLLaMA model Zhang et al. (2024), followed by supervised fine-tuning. More details on the experimental setup can be found in Appendix I.
@@
TinyLLaMA (VE) chat 22.61 ($+1.31$)
TinyLLaMA (PVE) chat 29.18 ($+7.88$)
@@
6.2 Compression Ratios
An encoding comparison was conducted on a consistent corpus to evaluate the compression efficiency of the vocabularies from LLaMA (AceGPT) and AraLLaMA, using LLaMA as the benchmark. AraLLaMA notably enhanced the baseline by achieving a token compression ratio of 0.3174, following the augmentation of its vocabulary with 12,800 Arabic subwords.
@@ 6.3 Benchmarking in English dataset
We evaluated the accuracy of both base and chat models on the English MMLU dataset. As illustrated in Table 10 (shown in Appendix F), in the base model category, AraLLaMA’s accuracy is slightly lower than that of the original LLaMA model but notably higher than the AceGPT model, which is also trained on the LLaMA architecture. This indicates that expanding Arabic capabilities via an annealing algorithm does not compromise the model’s inherent English proficiency. This offers a viable solution for language transfer in large models. After undergoing SFT, AraLLaMA achieves the highest accuracy among models of similar size and surpasses the Jais-30B model, which has a greater number of parameters.
@@ Appendix A Related Work
Recent efforts have centered on developing open-source LLMs as alternatives to proprietary models like GPT-3.5 Turbo and GPT-4 (Taori et al., 2023; Chiang et al., 2023; Conover et al., 2023; Chen et al., 2023; Sengupta et al., 2023). These initiatives have expanded beyond English, addressing languages with fewer available resources and creating models specifically tailored to diverse linguistic landscapes (Chen et al., 2023; Üstün et al., 2024). SeaLLMs (Nguyen et al., 2023) are adapted from English-centric models by extending vocabulary and fine-tuning to better capture regional language complexities. Jais (Sengupta et al., 2023) introduces a model trained from scratch based on GPT architecture, while AceGPT Huang et al., 2023a offers a model designed to adapt to local Arabic culture, specifically tailored to regional nuances. This trend highlights the growing need for multilingual LLMs that perform well in low-resource environments while maintaining competitive performance against more established models.
@@ Appendix B CEFR Language Proficiency Levels
Appendix C Arabic data distribution
Table 7 show the Arabic dataset primarily draws from several key sources, with the largest contribution coming from the Common Crawl (filtered) dataset, which accounts for 55.5% of the total data. Other significant sources include WebText, which contributes 26.7%, and Books+Newspapers, providing 8.9% with 2.5 billion tokens. Additionally, Wikipedia is divided into two parts, contributing 3.76% and 5.14%. These diverse sources collectively form the foundation for training the Arabic language model.
@@ Appendix C Arabic data distribution
Appendix D Data mixture
@@
Appendix E Comparison of compression ratio and OOV changes at different stages between exponential and uniform expansion
@@
Appendix F Evaluation of models in English MMLU dataset
@@
Model STEM Human- ities Social Sciences Others Avg. STEM Social Sciences Human- ities Other Avg. Avg.
@@
ChatGPT 3.5 Turbo 58.39 72.12 78.02 69.95 69.62 - - - - - -
Table 10: Evaluation of models in English MMLU dataset: few-shot on base model and zero-shot on chat model. Under the zero-shot setting, the LLaMA2-13B model does not follow instructions for unknow reason.
Appendix G ALAN examples

### v2

@@ 4.2 Training details
The refined methodology for LLaMA2 model’s vocabulary expansion incorporated 12,800 Arabic subwords derived through the I-BPE algorithm. The initialization procedure for each new token employed decomposition into constituent subwords from the original LLaMA2 vocabulary, with embedding initialization achieved via averaging the embeddings of these component tokens. This initialization strategy preserves semantic relationships between new and existing tokens, thereby enhancing training stability and facilitating vocabulary integration.
The training procedure was structured into 16 distinct stages  [^3] Although incremental token addition is theoretically feasible, a staged implementation ($N=16$) was chosen to simplify data preparation., each processing 30B tokens, culminating in a total corpus of 480B tokens. A cosine annealing schedule governed the proportion of Arabic to English content, with Arabic representation increasing systematically from 30% to 90% across stages. This progressive exposure enables gradual adaptation to Arabic linguistic structures while preserving cross-lingual transfer capabilities via continued English exposure. Mathematical and programming content was maintained at a consistent 5% throughout all stages to ensure robust inference capabilities in these domains (see Appendix E). The final training distribution comprised approximately 251.4B Arabic tokens and 204.6B English tokens.
The pre-training framework consisted of two principal epochs: an initial epoch employing vocabulary annealing for data distribution optimization, followed by a secondary epoch utilizing the fully refined vocabulary. An additional 20B tokens of training data were processed subsequent to vocabulary expansion to further enhance model performance. Each training phase implemented a discrete cosine learning rate schedule with warm-up period, producing a vocabulary-specific model at its conclusion, thereby rendering each phase functionally independent.
This stage-wise approach facilitates systematic integration of new tokens, enabling the model to adapt to evolving data representations while developing comprehensive understanding of linguistic patterns. The graduated modulation of language distribution—progressively increasing Arabic content while decreasing English representation—optimizes the model’s capacity to process Arabic while maintaining cross-lingual capabilities.
The implementation utilized LLaMA2 architecture in 7B and 13B parameter configurations, trained on 2,368 Ascend 910A processors. Training durations were 7 and 11 days for the 7B and 13B models, respectively. The computational configuration employed model parallelism of degree 2 and pipeline parallelism of degree 4. Optimization was conducted using AdamW with 4,096-token context length. Each training stage utilized a cosine learning rate scheduler initialized at 1e-5 and decaying to 2e-6, with a 15% warm-up interval. Gradient accumulation factor 8 yielded an effective batch size of 4,736, enabling processing of approximately 0.019B tokens per batch.
@@ 5.1 Experimental settings
To assess world knowledge, we employ four widely used benchmarks. MMLU (Measuring Massive Multitask Language Understanding) Hendrycks et al. (2021a) evaluates knowledge acquired during pretraining across a broad range of subjects; we utilize both the original English version and the Arabic version introduced in Huang et al. (2024) to ensure multilingual coverage. RACE (Reading Comprehension from Examinations) serves as a large-scale English reading comprehension benchmark that focuses on educational knowledge. EXAMS (Multi-subject High School Examinations Dataset for Cross-lingual and Multilingual Question Answering) further expands coverage by including subject-diverse exam questions drawn from multiple languages and curricula. ArabicMMLU complements these by providing an Arabic-specific variant of MMLU, tailored to reflect regional knowledge across various Arab countries and subjects. Beyond general knowledge evaluation, we also examine cultural and value alignment using ACVA-all and ACVA-clean, which focus on Arabic cultural relevance and localization. To comprehensively evaluate inference and reasoning abilities, we translate two commonsense reasoning benchmarks of varying difficulty—BoolQ and ARC-Challenge (ARC-C)—into Arabic, allowing for consistent cross-lingual assessment.
To ensure a fair comparison of candidate models, we adhere to the settings established for each benchmark separately. Furthermore, for translated benchmarks, we utilize the generation approach evaluation method as outlined in Huang et al. (2024). Specifically, we employed GPT-3.5-Turbo-1106 to translate datasets from English to Arabic for benchmarks that were not originally in Arabic.
@@
To compare LLMs trained or available in Arabic, we have selected several prominent Arabic LLMs or multilingual LLMs as baselines for comparison: (1) AceGPT-[7B,13B]: This set includes fully fine-tuned generative text models based on LLaMA2, specifically customized for the Arabic domain. (2) Mistral-7B-Instruct-v0.2 Jiang et al. (2023): The fine-tuned model achieves a balance between performance and efficiency. (3) Jais-[13B,30B] Sengupta et al. (2023): A pre-trained bilingual large language model designed for both Arabic and English. (4) Bloom-[7B]: A multilingual language model extensively trained on diverse textual data, allowing it to produce fluent text in 46 languages and 13 programming languages. (5) LLaMA2-[7B,13B]: A popular and competitive baseline model in the general domain. (6) OpenAI GPT: This includes GPT4 and ChatGPT, closed-source LLMs also strong at multilingual tasks.
@@ 5.2 Evaluation Results
In our study, the performance of base models was assessed on two Arabic-specific MMLU datasets: Arabic MMLU translate Huang et al. (2024) and ArabicMMLU Koto et al. (2024). The left side of Table 12 details the models’ accuracies on the Arabic MMLU translate dataset within a few-shot setting. It is evident from the data that the AraLLaMA-7B-base and AraLLaMA-13B-base models exhibit superior accuracy rates compared to models of similar scale. Notably, the AraLLaMA-13B-base model outperforms the Jais-30B model, which has a significantly larger parameter count.
@@
Models Arabic-trans MMLU Huang et al. (2024) ArabicMMLU Koto et al. (2024) Total
STEM
Human-
ities
Social
Sciences
Others Avg. STEM
Social
Sciences
Human-
ities
Arabic
Language
Other Avg. Avg.
@@
GPT-3.5 Turbo 43.38 44.12 55.57 53.21 49.07 53.8 57.0 57.5 57.6 63.8 57.7 53.39
Table 2: Evaluation of base models. We adopt a few-shot setting on Arabic-translated MMLU Huang et al. (2024) and a zero-shot setting with option logit probability in ArabicMMLU Koto et al. (2024). Numbers with the best performance are in bold in 7B and 13B groups.
@@
MMLU
MMLU
Koto et al. (2024)
EXAMS
ACVA
clean
ACVA
all
BoolQ
ARC-C
Avg. BoolQ RACE Avg. Avg.
@@
Table 3 presents the comprehensive evaluation results across various benchmarks for the candidate models, spanning from Arabic to English. Overall, AraLLaMA outperforms all baseline models in the Arabic tasks. Particularly noteworthy is its proficiency in knowledge-related evaluations such as Arabic-translated MMLU and EXAMS, surpassing other models by at least 1.3%. This highlights the model’s expertise in addressing Arabic knowledge-related questions. Additionally, AraLLaMA demonstrates strong performance in tasks related to Arabic culture and value alignment. In terms of commonsense reasoning, AraLLaMA exhibits notable skills in tasks such as the translated versions of BoolQ and ARC-Challenge, showcasing its reasoning capabilities in Arabic. Beyond Arabic benchmarks, we also investigated the English proficiency of the models to determine whether specialization in one language affects performance in the other. The results indicate that the model maintains its English proficiency and displays robustness in multilingual assessments. It is noteworthy that the lower accuracy of the Jais is attributed to its refusal to answer for unknown reasons.
In a comprehensive evaluation of the ACVA dataset aimed at gauging the understanding of Arabic cultural nuances under a zero-shot setting, our AraLLaMA models showcased unparalleled performance. The AraLLaMA-13B-chat, in particular, stood out with exceptional Average F1 scores of 76.37% and 76.90% in “all set" and "clean Set" categories, respectively, even outperforming the renowned GPT-3.5 Turbo in the "All set" category. This performance not only highlights the AraLLaMA models’ superior grasp of Arabic culture but also establishes them as leading figures among open-source models in this nuanced domain. Compared to other top-tier open-source contenders, including the Jais-30B-chat variants, the AraLLaMA-13B-chat model’s superior results. The instruction-following tests can be found in Appendix I.
@@ 6 More Analysis
Model STEM
Social
Sciences
Humanities
Arabic
Language
Other Avg.
@@ 6.1 Ablation Study on Progressive Vocabulary Expansion
To further demonstrate the effectiveness of progressive vocabulary expansion in downstream task adaptation, we conduct continuous pre-training on a 1B-parameter TinyLLaMA model Zhang et al. (2024), followed by supervised fine-tuning. More details on the experimental setup can be found in Appendix J.
@@
TinyLLaMA (VE) chat 22.61 (+1.31+1.31)
TinyLLaMA (PVE) chat 29.18 (+7.88+7.88)
@@ 6.2 Benchmarking in English dataset
We evaluated the accuracy of both base and chat models on the English MMLU dataset. As illustrated in Table 12 (shown in Appendix G), in the base model category, AraLLaMA’s accuracy is slightly lower than that of the original LLaMA model but notably higher than the AceGPT model, which is also trained on the LLaMA architecture. This indicates that expanding Arabic capabilities via an annealing algorithm does not compromise the model’s inherent English proficiency. This offers a viable solution for language transfer in large models. After undergoing SFT, AraLLaMA achieves the highest accuracy among models of similar size and surpasses the Jais-30B model, which has a greater number of parameters.
6.3 Decoding Efficiency Analysis
We conducted a systematic evaluation of generation efficiency between LLaMA2 and AraLLaMA 7B chat models on Arabic text generation tasks. Each model was tested on standardized Arabic prompts with a maximum output length of 100 tokens. To ensure statistical reliability, we performed five independent trials and analyzed only Arabic language outputs, excluding any non-Arabic tokens from the performance calculations.
Model Tokens/Second Words/Second
LLaMA2 $29.68\pm 0.04$ $4.55\pm 0.50$
AraLLaMA $30.12\pm 0.06$ $20.37\pm 0.04$
Table 6: Comparative analysis of generation speed between LLaMA2 and AraLLaMA on Arabic text.
Table 6 shows that while both models achieve similar token processing speeds ( 30 tokens/second, $p>0.05$), AraLLaMA generates words 4.5× faster. This efficiency gain (from $4.55\pm 0.50$ to $20.37\pm 0.04$ words/second) demonstrates the effectiveness of our vocabulary expansion approach. The improved word-level performance while maintaining similar token-level speeds indicates that our language-specific tokenization strategy successfully optimizes text generation for Arabic’s morphological complexity.
@@ Appendix A Related Work
Recent efforts have centered on developing open-source LLMs as alternatives to proprietary models like GPT-3.5 Turbo and GPT-4 (Taori et al., 2023; Chiang et al., 2023; Conover et al., 2023; Chen et al., 2023; Sengupta et al., 2023). These initiatives have expanded beyond English, addressing languages with fewer available resources and creating models specifically tailored to diverse linguistic landscapes (Chen et al., 2023; Üstün et al., 2024). SeaLLMs (Nguyen et al., 2023) are adapted from English-centric models by extending vocabulary and fine-tuning to better capture regional language complexities. Jais (Sengupta et al., 2023) introduces a model trained from scratch based on GPT architecture, while AceGPT Huang et al., 2024 offers a model designed to adapt to local Arabic culture, specifically tailored to regional nuances. This trend highlights the growing need for multilingual LLMs that perform well in low-resource environments while maintaining competitive performance against more established models.
@@
Aspect Benchmark
Language
(+ translation)
Size Evaluation Types Metrics
Knowledge Ability
RACE Lai et al. (2017) EN 4.9K Multiple-choice Questions Accuracy
MMLU Hendrycks et al. (2021b) EN (+AR) 14K Multiple-choice Questions Accuracy
Arabic Cultural
and Value Alignment
ACVA-all Huang et al. (2024) AR 9K Yes/No binary Questions F1-score
ACVA-clean AR 2.48K Yes/No binary Questions F1-score
Commonsense
Reasoning
BoolQ Clark et al. (2019) EN (+AR) 3.27K Yes/No binary Questions Accuracy
Table 7: Overview of Evaluation benchmarks
@@ Appendix B CEFR Language Proficiency Levels
Appendix C Cognitive Mechanisms of Vocabulary Acquisition
We reviewed relevant literature to confirm the phenomenon of exponential vocabulary expansion in second language acquisition and the cognitive theories that support it. Studies indicate that learners typically begin by mastering a small set of high-frequency vocabulary in the early stages of language learning. As they progress, their vocabulary size grows rapidly. This process can be explained through the following two aspects:
Cognitive Mechanisms of Incremental Learning
In the initial stages, learners build their understanding by repeatedly encountering and using simple foundational words. Research by Krashen (1982) and Nation and Nation (2001) shows that mastering high-frequency vocabulary is crucial for understanding more complex linguistic structures. These foundational words provide a stable cognitive base, allowing learners to gradually expand their vocabulary Zhang et al. (2021); Nakata (2015).
Exponential Vocabulary Growth
Once learners acquire foundational vocabulary, the rate of vocabulary expansion accelerates. Through extensive reading and structured learning strategies such as spaced retrieval practice León Romero et al. (2016), learners are able to acquire complex vocabulary in a relatively short period. Coady (1996) emphasize that extensive reading provides a large amount of language input, enabling learners to incrementally encounter and absorb more advanced vocabulary.
Appendix D Arabic data distribution
Table 9 show the Arabic dataset primarily draws from several key sources, with the largest contribution coming from the Common Crawl (filtered) dataset, which accounts for 55.5% of the total data. Other significant sources include WebText, which contributes 26.7%, and Books+Newspapers, providing 8.9% with 2.5 billion tokens. Additionally, Wikipedia is divided into two parts, contributing 3.76% and 5.14%. These diverse sources collectively form the foundation for training the Arabic model.
@@ Appendix D Arabic data distribution
Appendix E Data mixture
@@
Appendix F Comparison of compression ratio and OOV changes at different stages between exponential and uniform expansion
@@
Appendix G Evaluation of models in English MMLU dataset
@@
Model STEM
Human-
ities
Social
Sciences
Others Avg. STEM
Social
Sciences
Human-
ities
Other Avg. Avg.
@@
GPT-3.5 Turbo 58.39 72.12 78.02 69.95 69.62 - - - - - -
Table 12: Evaluation of models in English MMLU dataset: few-shot on base model and zero-shot on chat model. Under the zero-shot setting, LLaMA2-13B model does not follow instructions for unknow reason.
Appendix H ALAN examples